The average speed of a body which covers half the distance with velocity v1 and remaining half distance with velocity v2 is
1
Expert's answer
2021-09-02T11:35:50-0400

The average speed is given by

"v_{\rm ave}=\frac{l_{tot}}{t_{tot}}=\frac{l/2+l/2}{l/2v_1+l/2v_2}=\frac{2v_1v_2}{v_1+v_2}"